mirror sites: PL (internal link) HQ     [?]

Health Check monitor

 

CAL | HC | refs | QC
HealthCheck Monitor
   HOME | UsersGuide
  sl icon ALL INSTRUMENTS
 
FULL reports
VIMOS:
detector: bias
Other HC:
QC links:
        [page auto-refreshes after 300 sec]   [stop | on]            [press Ctrl+Shift+R to enforce refresh of scores, dates and news]
VIMOS trending system: HEALTH CHECK report
HELP USERS-GUIDE MORE
Last update: 2014-04-19T13:42:56 (UT)
  | now: 2014-04-19T14:48:26 (UT) muc03 QC pipeline: vimos-2.9.13 (installed 2014-03-19)
same group: median_master ron_master rms_master ron_raw median_vs_temp   averages_low averages_high
General news:
VIMOS news:  
Report news:
DATE*: [?] 2014-04-12
report | NLT
2014-04-13
report | NLT
2014-04-14
report | NLT
2014-04-15
report | NLT
2014-04-16
report | NLT
2014-04-17
report | NLT
2014-04-18
report | NLT
P...
Product quality:
products products products products products products products
  last OPSLOG data:
2014-04-18*

last files ...
monitor

scores&comments | history ... | plot tutorial ... | contact                 daily/often; important to check     [?]
*Date on this monitor changes at 21:00 UT
Print: Information and research:
- png file - advanced studies: QC1 browser | QC1 plotter
Plot
?
Symb
?
Source
*
OPS?
**
Average ? Thresholds ? N_
data
QC1
parameter
Data
downloads
Remarks
method value unit method value
1 QC1DB yes (5) MEDIAN 177 ADU VAL170,182 91 median_master this | last_yr | all Q2, median bias, low gain
2 QC1DB yes (5) MEDIAN 200 ADU VAL195,207 91 median_master this | last_yr | all Q1, median bias, low gain
3 QC1DB yes (5) MEDIAN 194 ADU VAL160,220 91 median_master this | last_yr | all Q2, median bias, high gain
4 QC1DB yes (5) MEDIAN 195 ADU VAL190,217 91 median_master this | last_yr | all Q1, median bias, high gain
5 QC1DB yes (5) MEDIAN 202 ADU VAL196,207 91 median_master this | last_yr | all Q3, median bias, low gain
6 QC1DB yes (5) MEDIAN 204 ADU VAL200,210 91 median_master this | last_yr | all Q4, median bias, low gain
7 QC1DB yes (5) MEDIAN 205 ADU VAL195,210 91 median_master this | last_yr | all Q3, median bias, high gain
8 QC1DB yes (5) MEDIAN 200 ADU VAL195,210 91 median_master this | last_yr | all Q4, median bias, high gain
*Data sources: QC1DB: QC1 database; LOCAL: local text file | **OPS: OPSLOG data are included
Plot 1
scores: 8|
data source:vimos_bias
(QC1 database)
dataset:median_master
median:177ADU
fixed thresholds:170...182ADU
N_data plotted:91
[click on plot for closeup]
Plot 2
scores: 8|
data source:vimos_bias
(QC1 database)
dataset:median_master
median:200ADU
fixed thresholds:195...207ADU
N_data plotted:91
[click on plot for closeup]
Plot 3
scores: 9|
data source:vimos_bias
(QC1 database)
dataset:median_master
median:194ADU
fixed thresholds:160...220ADU
N_data plotted:91
[click on plot for closeup]
Plot 4
scores: 9|
data source:vimos_bias
(QC1 database)
dataset:median_master
median:195ADU
fixed thresholds:190...217ADU
N_data plotted:91
[click on plot for closeup]
Plot 5
scores: 8|
data source:vimos_bias
(QC1 database)
dataset:median_master
median:202ADU
fixed thresholds:196...207ADU
N_data plotted:91
[click on plot for closeup]
Plot 6
scores: 8|
data source:vimos_bias
(QC1 database)
dataset:median_master
median:204ADU
fixed thresholds:200...210ADU
N_data plotted:91
[click on plot for closeup]
Plot 7
scores: 9|
data source:vimos_bias
(QC1 database)
dataset:median_master
median:205ADU
fixed thresholds:195...210ADU
N_data plotted:91
[click on plot for closeup]
Plot 8
scores: 9|
data source:vimos_bias
(QC1 database)
dataset:median_master
median:200ADU
fixed thresholds:195...210ADU
N_data plotted:91
[click on plot for closeup]
This plot

Trending plots for bias QC parameters.

All bias parameters are measured in the central 1600x1800 pixels of the detectors. The median bias level and the rms of the master should show a stable trend without sudden jumps or increase/decrease. The read-out noise (ron_raw) is calculated from the difference of the first two raw frames. Some outliers with high values are occasionally observed due to cosmic ray events. The read-out noise as measured from master frames (ron_master) should be extremely stable.

The new detectors (since July 2010) show a correlation of the bias level with temperature, especially for quadrant 2.

Bias frames are usually measured daily during day-time. They come in two different varieties: with low gain (used for imaging) and with high gain (used to calibrate the spectroscopic modes).

Template IDs: img_cal_Dark and spec_cal_Dark. Pipeline recipe: vmbias.  


General information

Click on any of the plots to see a close-up version.

If applicable, the latest values from Paranal ops logs are plotted as blue dots. More information about those values, including a download option, is available in the OPS column.

The latest date is indicated on top of the plot, data points belonging to that date are specially marked.

If configured,

  • statistical averages are indicated by a solid line, and thresholds by broken lines
  • outliers are marked by a red asterisk. They are defined as data points outside the threshold lines
  • "aliens" (= data points outside the plot Y limits) are marked by a red arrow ( or )
  • you can download the data for each parameter set if the 'Data downloads' link shows up